OpenStudy (anonymous):

find the complex zeros of f(x)=x^3+125

OpenStudy (anonymous):

\[x^3+125=x^3+5^3=(x+5)(x^2-5x+25)\] Find the roots by setting this equal to 0.

OpenStudy (anonymous):

thank you